### 04 Jan 2024 v1

# BA-SAM: Scalable Bias-Mode Attention Mask for Segment Anything Model

Abstract
In this paper, we address the challenge of image resolution variation for the Segment Anything Model (SAM). SAM, known for its zero-shot generalizability, exhibits a performance degradation when faced with datasets with varying image sizes. Previous approaches tend to resize the image to a fixed size or adopt structure modifications, hindering the preservation of SAM’s rich prior knowledge. Besides, such task-specific tuning necessitates a complete retraining of the model, which is cost-expensive and unacceptable for deployment in the downstream tasks. In this paper, we reformulate this issue as a length extrapolation problem, where token sequence length varies while maintaining a consistent patch size for images of different sizes. To this end, we propose Scalable Bias-Mode Attention Mask (BA-SAM) to enhance SAM’s adaptability to varying image resolutions while eliminating the need for structure modifications. Firstly, we introduce a new scaling factor to ensure consistent magnitude in the attention layer’s dot product values when the token sequence length changes. Secondly, we present a bias-mode attention mask that allows each token to prioritize neighboring information, mitigating the impact of untrained distant information. Our BA-SAM demonstrates efficacy in two scenarios: zero-shot and fine-tuning. Extensive evaluation on diverse datasets, including DIS5K, DUTS, ISIC, COD10K, and COCO, reveals its ability to significantly mitigate performance degradation in the zero-shot setting and achieve state-of-the-art performance with minimal fine-tuning. Furthermore, we propose a generalized model and benchmark, showcasing BA-SAM’s generalizability across all four datasets simultaneously.
Figure 1: (a) Top: contrast between prior methods [79, 4] and BA-SAM. For large-scale datasets, previous approaches often resize images or change patch sizes to handle the issue of varying resolutions. In contrast, we propose a Scalable Bias-Mode Attention Mask (BA-SAM), which enhances SAM’s adaptability to varying image resolutions while eliminating structure modifications. Bottom (left): We introduce a generalized model that does not require any task-specific designs, yet outperforms state-of-the-art methods across four datasets. Bottom (right): With resolution variations, prior models’ performance degrades drastically. Instead, BA-SAM consistently alleviates this issue.
footnotetext: *Equal contribution.footnotetext: Corresponding author.
1 Introduction
Recently, the computer vision community [30, 66, 29, 9, 74, 26, 85, 84, 86, 83, 87, 88, 89, 31, 90, 22, 58, 50, 49] has experienced a surge in the development of various foundation models [54, 35, 21]. Notably, Meta has introduced SAM (Segment Anything Model) [38], a prompt model that has made a significant impact. SAM can segment any object in an image or video by incorporating a single visual prompt, such as a box or a point, without requiring additional training. SAM is trained on an extensive SA-1B dataset [38], consisting of over 11 million images and one billion masks. Its emergence has undeniably showcased robust generalization capabilities across diverse images and objects, paving the way for new possibilities and avenues in intelligent image analysis and understanding [76, 8, 79, 34]. Based on SAM, Some variants have been proposed, such as MobileSAM [76] and SAM-Adapter [8]. These efforts typically focus on improving SAM’s performance on specific datasets.
During the pre-training of SAM [38], the input image size is fixed at 1024. As a foundational model, SAM is expected to exhibit generalization capabilities across various downstream tasks, each associated with datasets featuring different image sizes. This is particularly crucial for high-resolution (HQ) datasets characterized by larger dimensions and more details. SAM performs well when the resolutions align with its training resolution of 1024. However, significant performance degradation is observed when inferring with resolutions larger than 1024. Hence, we aim to study a practical and realistic problem to enhance SAM’s adaptability to varying image resolutions of different datasets.
Since SAM adopts the standard Vision Transformer [14] architecture, there are two common approaches to address the inconsistency between training and inference sizes for the ViT architecture. As depicted in Figure 1, the first approach, e.g., MSA [79] and SAM-Adapter [8], involves directly resizing all datasets to match a predefined size. Conversely, the second approach, exemplified by FlexiViT [4], entails adjusting the patch size to accommodate larger image resolutions. Nevertheless, tuning the image or patch size necessitates a complete retraining of the model, which is cost-expensive and unacceptable for deployment in the downstream tasks. Besides, it prevents leveraging the rich prior knowledge reserved in the pre-trained model of SAM. As such, our objective is to explore a solution that enhances SAM’s adaptability to datasets of varying resolutions while avoiding structural modifications to SAM.
In this paper, we introduce a novel perspective that reframes the challenge of image resolution variation as a length extrapolation problem. Specifically, as depicted in Figure 1, for images of varying sizes, we employ different token sequence lengths while keeping a consistent patch size. It has been observed that the inconsistency in token length between training and prediction is a key factor in performance degradation. This inconsistency manifests in two aspects: Firstly, changes in token length lead to variations in the magnitude of attention module values. When the dot product result becomes significantly large in magnitude, it can drive the subsequent Softmax layer into regions with minimal gradients. Consequently, the attention distribution after Softmax becomes highly concentrated, giving rise to the issue of vanishing gradients. Secondly, longer predictions rely on untrained information, such as additional position encodings. The introduction of untrained parameters brings a considerable amount of noise to the model, which, in turn, affects its performance.
To address these issues, we propose a Scalable Bias-Mode Attention Mask (BA-SAM) to enhance the length extrapolation capability of SAM. Our approach introduces two novel designs. Firstly, to ensure consistency in the attention layer’s dot product value, we present an improved scaling factor. This factor effectively regulates the magnitude of values within the attention layer, mitigating disruptive effects resulting from substantial changes in dot product operations and context length. Secondly, with a focus on maintaining consistency in attention focus areas, we introduce a novel bias-mode attention mask. This attention mask penalizes attention scores between distant query-key pairs, with the penalty increasing as the distance between the key and query grows. Consequently, when the context length varies, the influence of untrained distant information on each token diminishes. We achieve this mask by adding a bias after the query-key dot product, and this design is highly lightweight and could be seamlessly integrated into SAM-based models with minimal computational overhead.
Our approach demonstrates efficacy in two scenarios: zero-shot and fine-tuning. Extensive evaluation on datasets from five diverse tasks are conducted, including DIS5K [53], DUTS [64], ISIC [12], COD10K [16], and COCO [45]. These datasets vary in resolution, mostly exceeding SAM’s default resolution of 1024. In the zero-shot setting, our BA-SAM alleviates the model’s performance degradation caused by expanding the inference resolution without requiring any additional training. With a few fine-tuning epochs on downstream tasks, our BA-SAM consistently achieves state-of-the-art accuracy across all datasets. Additionally, to further demonstrate BA-SAM’s generalizability, we propose a generalized model and a new benchmark, which utilize one model to attain state-of-the-art performance across all four datasets simultaneously.
2 Related Work
Visual Foundation Models. Models that are trained on broad and can be adapted to numerous downstream tasks are called “Foundation Models” [5, 67, 43]. These models, Vision-Language Models (VLM) (CLIP [54] and DALL-E [55]) combine computer vision and natural language processing to understand and generate descriptions or analyze visual content using textual and visual information. Masked Image Modeling [73, 48] (MIM) refers to a technique where parts of an image are masked during training to encourage a model to learn contextual information and complete missing regions. SAM [38] is a model designed for segmenting objects or areas in images, offering precise segmentation capabilities. In this paper, we use a variant of SAM called MobileSAM [76] as the baseline method. As our design optimizes the Transformer itself, it applies to all Transformer-based vision foundation models.
Figure 2: Illustration of the proposed BA-SAM method. (a) In the original SAM, when input token sequences length varies during testing, the magnitude of the Softmax outputs changes drastically. In contrast, we propose a new scaling factor to address this issue. (b) We introduce a bias-mode attention mask, which produces increasing penalties on attention scores as the distance between query and key grows.
Resolution Variation Processing. To enable models to be more adaptable to variations in resolutions, traditional methods to deal with VIT have relied on adjustments to positional embeddings [40] and patch sizes [4, 6, 24, 75, 44, 30, 39]. Patch n’ Pack [13] employed sequence packing during training to handle inputs with arbitrary resolutions and aspect ratios. All of them necessitate training from scratch, incurring substantial computational and time costs. In contrast to previous approaches, we extend the concept of ’length extrapolation’ from NLP into the context of addressing scale variations in CV. Length extrapolation refers to a model’s ability to generalize well to longer inputs than those it was trained on. In NLP, it has been successfully used, such as in ALIBI [52] and KERPLE [11], to enable models to adapt to longer sequences without significant performance degradation. Our approach seamlessly extends to two scenarios: zero-shot and finetuning, allowing us to leverage prior knowledge embedded in the SAM and significantly reduce training efforts.
Parameter Efficient Tuning. There have been some pioneering works for the Parameter Efficient Tuning (PETL) of visual models, such as AdaptFormer [7] and visual prompt tuning (VPT) [36]. He et al. [28] analyzed the unified view among PETL techniques such as prefixtuning [42], Prompt-tuning [36], and adapter [7]. Our method belongs to the category of Parameter Efficient Tuning.
Visual Attention Modeling. Various studies have incorporated attention mechanisms into neural network architectures designed for visual tasks [63, 68, 41, 3, 33, 93]. These mechanisms are employed in a channel-wise manner to capture cross-feature information [66, 10, 80]. They are also used for selecting paths in different branches of a network [61], or a combination of both strategies [78]. For instance, squeeze-and-excite network [32] features an attention-like module for modeling channel-wise relationships within layer features. Li et al. [41] utilize the attention mechanism to adapt the receptive field of neurons between network branches. The advent of transformers has led to hybrid architectures that integrate other modules. Bello’s work [2] introduces approximate content attention with a positional attention component. Child et al. [69] observe that many early layers in the network learn locally connected patterns akin to convolutions, indicating that hybrid architectures inspired by both transformers and convolutional networks are a compelling design choice. Several recent studies explore this approach for various tasks [59, 70, 65, 25]. In contrast to prior work, we do not introduce a new attention structure. Instead, we offer theoretical proof for optimizing existing attention mechanisms. This resulting optimization approach is applicable across various attention designs and demonstrates strong performance across multiple datasets.
3 Preliminaries
SAM. Segment Anything Model (SAM) [38] consists of three core modules: image encoder, prompt encoder, and mask decoder. The SAM model has been trained on the extensive SA-1B dataset [38], which comprises more than 1 billion automatically generated masks. This dataset contains 400 times more masks than any existing segmentation dataset (COCO [45], LVIS [27] and ADE20K [82]) and is accompanied by 11 million images and 1 billion masks. Consequently, SAM exhibits valuable and robust zero-shot generalization to new data without necessitating further training. For further details regarding the SAM methodology, readers are encouraged to refer to [38]. Our Scalable Bias-Mode Attention Mask (BA-SAM) focuses its optimization on the image encoder while keeping the structures of the mask decoder and prompt encoder unchanged.
Scaling Factor in Attention. The two most commonly used attention functions are additive attention [1] and dot-product attention [62]. The vanilla Transformer chooses dot-product attention because it is much faster and more space-efficient in practice. However, for large values of the dimension of the tokens $d_{k}$, the dot products grow large in magnitude, pushing the softmax function into regions with minimal gradients. They use the scaling factor $\lambda_{d}=\frac{1}{\sqrt{d_{k}}}$ to scale the dot products. To better analyze the role of the scaling factor, we express the output element $\mathbf{O}_{i}$ and the weight coefficient $a_{i,j}$ as follows:
$\mathbf{O}_{i}$ $=\sum_{j=1}^{N}a_{i,j}x_{j}v_{j},\quad a_{i,j}=\frac{e^{\lambda_{d}q_{i}\cdot k_{j}}}{\sum_{j=1}^{N}e^{\lambda_{d}q_{i}\cdot k_{j}}},$ (1)
where $\lambda_{d}$ represents the scaling factor.
Absolute & Relative Position Encoding in Attention. The original Transformer [62] incorporates absolute non-parametric positional encoding $p=(p_{1},\dots,p_{n})$ with the input token embedding $x$ as $x_{i}=x_{i}+p_{i}$, where $p_{i},x_{i}\in\mathrm{R^{d_{x}}}$. Other works replace them with parametric encoding [23] or adopted Fourier-based kernelized versions [51]. Absolute position encoding enforces a fixed size for all inputs. Recent work [56] considers the pairwise relationships between elements. Relative relation is important for tasks where the relative ordering or distance of the elements matters. This method encodes the relative position between input elements $x_{i}$ and $x_{j}$ into vectors $p_{i,j}^{v},p_{i,j}^{q},p_{i,j}^{k}\in\mathrm{R}^{d_{z}}$. Then, we reformulate Eq. (1) as follows:
$\mathbf{O}_{i}=\sum_{j=1}^{n}a_{i,j}\left(x_{j}v_{j}+p_{i,j}^{v}\right),$ (2)
$a_{i,j}=\frac{e^{\lambda\left(q_{i}+p_{i,j}^{q}\right)\cdot\left(k_{j}+p_{i,j}^{k}\right)}}{\sum_{j=1}^{n}e^{\lambda\left(q_{i}+p_{i,j}^{q}\right)\cdot\left(k_{j}+p_{i,j}^{k}\right)}},$ (3)
where $p_{i,j}^{v},p_{i,j}^{q},p_{i,j}^{k}$ is learned during training.
4 Methodology
From the above section, it is evident that the original SAM’s design has a substantial limit to handling the length extrapolation problem. To address this, as shown in Figure 2, we present a Scalable Bias-mode Attention Mask (BA-SAM). In Sec. 4.1, we provide a theoretical explanation for the scaling factor used in the original Transformer and introduce a new scaling factor to regulate the magnitude inconsistency caused by length extrapolation. In Sec. 4.2 we design a bias-mode attention mask to place more focus on neighboring tokens, ensuring the sensitivity of the model when meeting positional encoding variations. Finally, we explain how we will embed our proposed BA-SAM into the SAM-based structure in Sec. 4.3.
4.1 New Scaling Factor
We discovered from the original attention module in SAM [62] that when the dot product becomes significantly large in magnitude, it can drive the Softmax layer into regions with minimal gradients. This is because the attention distribution after Softmax becomes highly concentrated, giving rise to the issue of vanishing gradients. Upon closer examination of Eq. (1), it is obvious that the computation of the $q\cdot k$ term is intrinsically tied to both the token sequences length $N$ and the dimension $d_{k}$. When the token sequence length $N$ and/or the dimension $d_{k}$ significantly increases, the overall efficacy of the attention is impacted, thus leading to a noticeable performance reduction.
To address this, we attempt to design a new scaling factor that allows the model to cope with significant variations in $N$ and $d_{k}$ values. When $N$ or $d_{k}$ grows significantly, we expect to regulate the magnitude of the values within the attention layer, maintaining a similar magnitude. While [62] introduced a scaling factor $\lambda=\frac{1}{\sqrt{d_{k}}}$ to counteract the effect of the large growth in magnitude due to the dot products, the authors did not explain how this scaling factor was designed in the original paper. Below we will provide a theoretical derivation of this scaling factor, and then elaborate on our proposed new scaling factor.
The dimension $d_{k}$. We assume that the components of $q$ and $k$ are independent Gaussian distribution variables with mean 0 and variance 1, then we want to prove that the $q\cdot k$ has mean 0 and variance $d_{k}$. The mean of $q\cdot k$ is:
$\mathrm{E}[q\cdot k]$ $=\mathrm{E}\left[\sum_{i=1}^{d_{k}}q_{i}k_{i}\right]=\sum_{i=1}^{d_{k}}\mathrm{E}\left[q_{i}\right]\mathrm{E}\left[k_{i}\right]=0$ (4)
Similarly, we formulate the variance of $q\cdot k$ as follows:
$\operatorname{var}[q\cdot k]$ $=\operatorname{var}\left[\sum_{i=1}^{d_{k}}q_{i}k_{i}\right]=\sum_{i=1}^{d_{k}}\operatorname{var}\left[q_{i}\right]\operatorname{var}\left[k_{i}\right]=d_{k}$ (5)
Given this, we can approximately consider the $q\cdot k$ values to be within the range of $-3\sqrt{d_{k}}$ to $3\sqrt{d_{k}}$, according to properties of Gaussian distribution. For larger models, $d_{k}$ is generally a larger positive value, resulting in a significant increase in the magnitude of numerical values of $q\cdot k$, compared to the additive attention option which has the range of $\left[-3,3\right]$. Consequently, the attention distribution after Softmax becomes highly concentrated. This leads to severe gradient vanishing, which in turn hampers the effectiveness of training and can lead to less desired performance. As the $q\cdot k$ values lie in the range of $[-3\sqrt{d_{k}},3\sqrt{d_{k}}]$, the scaling factor can be simply defined as $\lambda_{d}=\frac{1}{\sqrt{d_{k}}}$, in order to maintain a similar magnitude.
Our new scaling factor. We provided the interpretation on how the original scaling factor was designed, which has been overlooked by the original authors. Now we explain the design of our new scaling factor.
Our goal is to regulate the magnitude of the Softmax value $a_{ij}$ (Eq. (1)) as the token sequences length $N$ and the dimension $d_{k}$ change. We further denote the token sequences length during the training as $N_{train}$ and the token sequence length during the testing as $N_{test}$, where $N_{test}>>N_{train}$. We now consider how to maintain a similar magnitude of $a_{ij}$ during training and testing times. We introduce the scaling factor $\lambda_{N}$, which is related to $N$ and controls the output scale of the Softmax value $a_{ij}$ , namely, $S(a_{i,j}^{N_{test}})=S(a_{i,j}^{N_{train}})$. Here, $S()$ represents the scale estimation function.
We next substitute Eq. (1) into $S(a_{i,j}^{N_{test}})=S(a_{i,j}^{N_{train}})$, and simplify $q_{i}\cdot k_{j}$ into $x_{i,j}$ provided the magnitude of $q\cdot k$ has been already maintained. Now we have:
$S(\frac{e^{\lambda_{n}x_{i,j}}}{\sum_{j=1}^{n}e^{\lambda_{n}x_{i,j}}})=S(\frac{e^{x_{i,j}}}{\sum_{j=1}^{m}e^{x_{i,j}}})$ (6)
$\Rightarrow S({e^{\lambda_{n}}})=S(\frac{N_{test}}{N_{train}})\Rightarrow\lambda_{n}=log_{N_{train}}N_{test}$
Considering both $\lambda_{d}$ and $\lambda_{n}$, we can ultimately derive our new scaling factor as:
$\lambda=\lambda_{d}\lambda_{n}=\frac{log_{N_{train}}N_{test}}{\sqrt{d_{k}}}$ (7)
Our new scaling factor in Eq. (7) ensures the magnitude of values during attention computation remains consistent, regardless of variations in $d_{k}$ and $N$. It will enhance the extrapolative capacity of the model.
4.2 Bias-Mode Attention Mask
Another challenge is that changes in token sequence length will lead to variations in positional encoding. It is important to ensure the insensitivity of the model when such positional encoding variations occur during testing.
One possible way is absolute encoding without trainable parameters, such as Sinusoidal [62]. It requires the position encoding to have strong local-to-global inference capabilities. As mentioned by Su [60], this resembles Taylor series approximations: by knowing the values of several derivatives at a particular point, effective estimates can be made within a neighborhood. Nevertheless, this assumes that the given function has high-order smoothness (higher-order derivatives exist and are bounded). Commonly used positional encodings are often combined with trigonometric functions, essentially high-frequency oscillatory functions with periodicity. However, these methods fail to satisfy the requirement of bounded high-order derivatives, making it less accurate to estimate the extrapolated results.
Another potential approach is using local attention, which constrains the model’s field of view and remains insensitive to variations in token sequence length. However, local attention is typically implemented using a local window, necessitating modifications to the SAM structure, which requires re-training from scratch and is unacceptable for deployment in the downstream tasks.
To this end, we propose to enable the attention layer to put more focus on neighboring tokens of the current token. As such, even with an increase in the length of a token sequence, each token is scarcely affected by the untrained tokens from distant positions. In particular, we design a simple yet effective bias-mode mask. We introduce a bias after the query-key dot product.
As shown in Figure 3, this mask exhibits a bias specified on the distance between the query-key pairs (i.e., $q\cdot k$). We expect that this proposed mask imposes penalties on attention scores between distant query-key pairs, and the penalty increases as the distance between a key $q$ and a query $k$ grows. To this end, we simply define the bias as the form of $b_{i,j}=\beta|i-j|$.
$a_{i,j}=\frac{e^{\lambda(q_{i}\cdot k_{j}+b_{i,j})}}{\sum_{j=1}^{n}e^{\lambda(q_{i}\cdot k_{j}+b_{i,j})}},$ (8)
where $\beta$ is a head-specific slope.
We further discuss the setting of $\beta$ based on different cases. When conducting zero-shot transfer learning without fine-tuning, we set $\beta$ to a static, non-learned fixed value. The specific value setting will be discussed in the experimental section (Sec. 5). When fine-tuning is required, we make $\beta$ trainable. Since our Bias-Mode Attention Mask is lightweight relative to the model structure, it incurs negligible training time overhead.
Figure 3: Embedding of our BA-SAM into a SAM backbone. NSF indicates our new scaling factor, and BM-AM denotes our designed bias-mode attention mask.
Figure 4: Visualization results of our BA-SAM on four object segmentation tasks, i.e., skin lesion segmentation, salient object segmentation, complex object segmentation, camouflaged object detection, which corresponds to four datasets: ISIC [12], DUTS [64], DIS-TE4 [53], and COD10K [16]. Our BA-SAM can handle the issue of varying image resolutions and segments accurately in different tasks.
4.3 BA-SAM Model
As shown in Figure 3, our BA-SAM is simple to implement, and can be seamlessly integrated into SAM [38] and its variants. Specifically, our design involves a new scaling factor (NSF) for the attention layer and a bias-mode attention mask (BM-AM). Our method does not involve any alterations to the model structure and is suited to both fine-tuning and non-fine-tuning cases. As for the fine-tuning case, it introduces negligible computational overhead as BM-AM incurs a very small amount of computation.
5 Experiments
Method
Train
Size
Test
Size
ISIC [12] $\Delta$diff DUTS [64] $\Delta$diff DIS-TE4 [53] $\Delta$diff COD10K [16] $\Delta$diff
Without fine-tuning
SAM [38] - 1024 0.421 - 0.298 - 0.362 - 0.217 -
- 2048 0.601 18.0% 0.360 6.2% 0.411 4.9% 0.391 17.4%
Ours (w [38]) - 1024 0.417 - 0.294 - 0.356 - 0.208 -
- 2048 0.589 17.2% 0.348 5.4% 0.406 5.0% 0.387 17.9%
MobileSAM [76] - 1024 0.463 - 0.502 - 0.544 - 0.465 -
- 2048 0.641 17.8% 0.437 6.5% 0.427 11.7% 0.346 11.9%
4096 0.693 23.0% 0.328 17.4% 0.355 18.9% 0.300 16.5%
Ours (w [76]) - 1024 0.452 - 0.486 - 0.515 - 0.440 -
- 2048 0.611 15.9% 0.413 7.3% 0.406 10.9% 0.321 11.9%
4096 0.657 20.5% 0.283 20.3% 0.361 15.4% 0.246 19.4%
With fine-tuning
MobileSAM [76] 1024 1024 0.098 - 0.034 - 0.057 - 0.030 -
2048 0.520 42.2% 0.082 4.8% 0.065 0.8% 0.063 3.3%
4096 0.542 44.4% 0.235 20.1% 0.101 4.4% 0.138 10.8%
2048 2048 0.083 - 0.045 - 0.056 - 0.036 -
4096 0.227 14.4% 0.091 4.6% 0.066 1.0% 0.059 2.3%
Ours (w  [76]) 1024 1024 0.037 - 0.031 - 0.055 - 0.029 -
2048 0.499 40.4% 0.078 4.4% 0.061 0.6% 0.057 2.5%
4096 0.531 43.6% 0.226 19.2% 0.098 4.3% 0.118 8.6%
2048 2048 0.080 - 0.043 - 0.053 - 0.033 -
4096 0.214 13.4% 0.088 4.4% 0.061 0.8% 0.056 2.3%
Table 1: Performance comparisons in varying image resolutions. We employed the widely-used MAE (Mean Absolute Error) score. Lower MAE scores indicate better model accuracy. $\Delta$diff denotes the performance degradation due to resolution changes. Compared to the SAM [38] and MobileSAM [76] baselines, our BA-SAM achieves smaller degradation when encountering token sequences length changes. The best MAE performance is highlighted in bold, and the smallest performance degradation is underlined.
We first describe the datasets and implementation details in Sec. 5.1. Then, we compare BA-SAM’s performance with state-of-the-art methods across multiple tasks in Sec. 5.2. Finally, extensive ablation studies and analysis are conducted in Sec. 5.3. to demonstrate the effectiveness.
5.1 Datasets and Implementations
Datasets. For a comprehensive evaluation of BA-SAM, we conduct extensive experiments on a wide range of segmentation tasks, i.e., salient object segmentation [18, 15, 20], complex object segmentation [19], skin lesion segmentation [93], camouflaged object detection [17], which correspond to four datasets: DUTS [64], DIS-TE4 [53], ISIC [12] and COD10K [16]. Besides, we also verify it on the challenging COCO [45] instance segmentation benchmark. More details about datasets can be referred to in the supplementary material.
Implementation Details. In zero-shot settings that do not require fine-tuning, we use the original SAM [38] backbone. For fine-tuning scenarios, we employ MobileSAM [76] as the backbone. MobileSAM [76] is a SAM variant with a structure similar to ViT-Tiny [72], and further details can be found in [76]. MobileSAM [76] uses a ViT-H-based SAM as a teacher network for distillation, ultimately achieving competitive accuracy compared to the original SAM but with significantly fewer parameters. For various object segmentation tasks, a random point is extracted from the ground truth as the prompt input during the fine-tuning phase. For instance segmentation, we use the ViT-B [14] backbone and the state-of-the-art detector Deformable-DETR  [91] trained on the COCO [45] dataset with Swin-L [47] backbone as box prompt generator. More details are provided in the supplementary material. The code will be released at this link.
Evaluation metrics. In the experiments, we use the widely used Mean Absolute Error (MAE) and Average Precision (AP) for evaluation. A lower MAE score and a higher AP score indicate better model performance.
5.2 Results
Results of Various Object Segmentation Tasks: Table 1 demonstrates the effectiveness of our approach across four diverse segmentation datasets. $\Delta$diff denotes the value of the performance degradation due to resolution changes during the inference. The upper and lower parts of the table indicate the results without and with fine-tuning, respectively due to input resolution variations. The best MAE performance is highlighted in bold, and the smallest degradation is underlined. We have three observations: Firstly, our proposed BA-SAM consistently outperforms both SAM [38] and MobileSAM [76] baselines on all four datasets. This is mainly because these baselines do not consider the issue of varying image resolutions. In contrast, our presented scaling factor and bias-mode attention mask explicitly handle this issue and further alleviate the performance degradation. Secondly, when testing on higher resolutions than the training size, SAM [38] and MobileSAM [76] baselines show less desirable results than the original image size. In contrast, our BA-SAM incurs significantly less performance drop in different datasets. Thirdly, during the experiment, we observe negligible computational overhead, whether fine-tuning is applied or not, which supports the claim in the method section. See Sec. 5.3 for details.
Results of Instance Segmentation: In Table 3, we evaluate the performance of our method on the COCO [45] instance segmentation benchmark. For a fair comparison, all experiments are conducted in a zero-shot manner, with the same initialization parameters for the comparative methods and without the use of any additional training data. Our BA-SAM consistently outperforms SAM [38] and MobileSAM [76] baselines, demonstrating better zero-shot generalization capability on instance segmentation.
Comparisons with State-of-the-Art Methods: To further demonstrate the superiority of the effectiveness and generalizability of our method, we compare the state-of-the-art approaches in Table 2. From the table, we have two following observations: Firstly, all the state-of-the-art approaches [57, 77, 46, 92, 81, 53, 37, 16] show less-desirable performances in each dataset. Instead, our BA-SAM (specialized models) consistently outperforms these methods when fine-tuned on each downstream dataset. Secondly, almost all of these state-of-the-art techniques are specifically designed for one task and cannot be generalized well to other tasks. Due to the strong zero-shot generalization capability of SAM [38], our proposed BA-SAM can also be employed as a generalized model, which fine-tunes with all of these downstream datasets in a unified and shared model. Importantly, unlike [46, 57, 77, 12], we eliminate the need for employing additional techniques to further enhance the performance. As shown in Table 2, our generalized model also consistently promotes the performance of SAM on all datasets, demonstrating its remarkable generalizability.
5.3 Ablation Study and Analysis
In this section, we first conduct ablation studies to study the contribution of each component. Then, we investigate the impact of the new scaling factor (NSF) and bias-mode attention mask (BM-AM) with a more detailed analysis.
Ablations Studies of Each Component. Table 4 summarizes the effect of each designed component on the settings with and without fine-tuning, respectively. The baseline means using the MobileSAM [76] as the base network that uses the vanilla scaling factor (VSF) in the attention layer [62]. New Scaling Factor and Bias-Mode Attention Mask are abbreviated as NSF and BM-AM, respectively. From the table, we observe that NSF could achieve superior performances than the baseline with VSF. This is because the vanilla attention in SAM [38] and MobileSAM [76] does not consider maintaining the magnitude consistency of the Softmax outputs changes drastically when input resolutions vary during testing. In contrast, our NSF explicitly maintains the magnitude consistency and alleviates the performance degradation. Furthermore, by adding BM-AM, the performance could be further boosted when extrapolating to a larger test length. These improvements confirm that these individual components are complementary and together they significantly promote the performance.
Methods ISIC [12] DUTS [64] DIS-TE4 [53] COD10K [16]
Specialized models
CASS [57] 0.086 - - -
DINO [77] 0.081 - - -
MSA [71] 0.049 - - -
[…]
IS-Net [53] - - 0.072 -
SINet [16] - - - 0.092
[…]
the same framework, 4 specialized models
Ours (BA-SAM) 0.037 0.031 0.055 0.029
generalized model
SAM [38] 0.419 0.298 0.373 0.217
Ours (BA-SAM) 0.054 0.030 0.054 0.054
Table 2: Comparison results (MAE) with state-of-the-art specialized models on various segmentation tasks.
Model AP $AP_{50}$ $AP_{75}$ $AP_{S}$ $AP_{M}$ $AP_{L}$
SAM [38] 42.5 69.6 44.7 29.7 47.0 56.7
Ours (w [38]) 43.0 70.0 45.4 30.0 47.6 57.4
MobileSAM [76] 40.8 68.4 41.6 26.0 44.4 57.6
Ours (w [76]) 41.6 69.0 42.3 26.2 44.8 58.4
Table 3: Results (AP) on COCO [45] instance segmentation.
Impact of Slope in Bias-Mode Attention Mask. In the Bias-Mode Attention Mask, the magnitude of the slope $\beta$ determines penalty rates in different heads. In the zero-shot case, we use a fixed slope. Table 5 reveals the impact of different slopes of BM-AM on the ISIC [12] dataset. We found that the best performance is achieved when $\beta=0.1$. Besides, our method is robust to different slope choices. In addition to the zero-shot setting, we also validate the BM-AM in fine-tuning scenarios on four datasets, shown in Table 1. The learnable parameters brought more performance growth compared with the fixed slope, while also bringing a slight increase in computation.
Computational Efficiency. In Table 6, we analyze the computational efficiency between the baselines and our BA-SAM. All the experiments are conducted on the same NVIDIA RTX 4090GPU to ensure fair comparisons. From the table, we observe that our BA-SAM is highly lightweight, incurring negligible computational overhead for the models. The reasons lie in two aspects: firstly, the NSF exhibits nearly identical computational complexity with the vanilla one. Besides, the BM-AM is seamlessly incorporated by adding a mask matrix to the query-key dot product before applying the softmax operation. Although there is a slight increase in memory usage, it remains negligible compared to the memory occupied by large models.
Methods ISIC [12] DUTS [64] DIS-TE4 [53] COD10K [16]
Without fine-tuning
Baseline [76] 17.8 6.5 11.7 11.9
+ NSF 16.4 7.5 11.3 12.0
+ BM-AM 16.8 7.2 11.7 11.9
+ Both 15.9 7.9 10.9 11.9
With fine-tuning (1024)
Baseline [76] 42.2 4.8 0.8 3.3
+ NSF 40.9 4.6 0.7 3.0
+ BM-AM 41.2 4.5 0.8 2.7
+ Both 40.4 4.4 0.6 2.5
With fine-tuning (2048)
Baseline [76] 14.4 4.6 1.0 2.3
+ NSF 13.1 4.6 0.8 2.4
+ BM-AM 13.7 4.5 0.9 2.3
+ Both 13.4 4.4 0.8 2.3
Table 4: Ablation study of each component on the settings with and without fine-tuning. Numbers indicate the performance degradation, $\Delta$diff. A lower $\Delta$diff means a better performance.
Size 0.1 0.5 1 1.5 2 2.5
MAE 0.0375 0.0383 0.0381 0.0393 0.0388 0.0391
$\Delta$diff 40.4 40.8 40.4 40.9 41.2 40.7
Table 5: Impact of the slope $\beta$ in BM-AM (zero-shot setting).
Model Params (M) Speed (ms) Train Hours (h)
SAM [38] 81 113.9 -
Ours (w [38]) 81 114.0 -
MobileSAM [76] 9.66 16.2 0.64
Ours (w [76]) 9.67 16.5 0.65
Table 6: Comparisons of computational efficiency between the baselines and our BA-SAM. Params: number of parameters. Speed: inference speed. The top part is performed on the zero-shot setting, and the bottom part is the scenarios with fine-tuning.
6 Conclusion
In this paper, we address the important problem of varying image resolutions in SAM models by reformulating it as a problem of length extrapolation. To enhance the length extrapolation capability of SAM, we propose Scalable Bias-mode Attention Mask (BA-SAM). A new scaling factor (NSF) is introduced to maintain the consistent magnitude of attention. In addition, a bias-mode attention mask (BM-AM) is designed to prioritize neighboring information, mitigating the impact of untrained distant information. Extensive evaluation on diverse datasets reveals its ability to significantly alleviate performance degradation in the zero-shot setting and achieve state-of-the-art performance with minimal fine-tuning. Furthermore, we proposed a generalized model and benchmark, showcasing BA-SAM’s generalizability across all four datasets simultaneously.

### 19 Mar 2024 v3

@@
 Furthermore, we propose a generalized model and benchmark, showcasing BA-SAM’s generalizability across all four datasets simultaneously.
-Figure 1: (a) Top: contrast between prior methods [79, 4] and BA-SAM.
+Figure 1: Top: contrast between prior methods [81, 4] and BA-SAM.
 For large-scale datasets, previous approaches often resize images or change patch sizes to handle the issue of varying resolutions.
 In contrast, we propose a Scalable Bias-Mode Attention Mask (BA-SAM), which enhances SAM’s adaptability to varying image resolutions while eliminating structure modifications.
-Bottom (left): We introduce a generalized model that does not require any task-specific designs, yet outperforms state-of-the-art methods across four datasets.
+Bottom (left): We introduce a generalized model that outperforms state-of-the-art methods across four datasets.
 Bottom (right): With resolution variations, prior models’ performance degrades drastically.
-Instead, BA-SAM consistently alleviates this issue.
+Instead, BA-SAM consistently alleviates this issue (The evaluation metric is MAE).
 footnotetext: *Equal contribution.footnotetext: Corresponding author.
 1 Introduction
-Recently, the computer vision community [30, 66, 29, 9, 74, 26, 85, 84, 86, 83, 87, 88, 89, 31, 90, 22, 58, 50, 49] has experienced a surge in the development of various foundation models [54, 35, 21].
+Recently, the computer vision community [29, 66, 28, 9, 75, 26, 87, 86, 88, 85, 89, 90, 91, 30, 92, 22, 58, 50, 49, 93] has experienced a surge in the development of various foundation models [54, 33, 21].
 Notably, Meta has introduced SAM (Segment Anything Model) [36], a prompt model that has made a significant impact.
@@
 Since SAM adopts the standard Vision Transformer [14] architecture, there are two common approaches to address the inconsistency between training and inference sizes for the ViT architecture.
-As depicted in Figure 1, the first approach, e.g., MSA [79] and SAM-Adapter [8], involves directly resizing all datasets to match a predefined size.
+As depicted in Fig.
+1, the first approach, e.g., MSA [81] and SAM-Adapter [8], involves directly resizing all datasets to match a predefined size.
 Conversely, the second approach, exemplified by FlexiViT [4], entails adjusting the patch size to accommodate larger image resolutions.
@@
 In this paper, we introduce a novel perspective that reframes the challenge of image resolution variation as a length extrapolation problem.
-Specifically, as depicted in Figure 1, for images of varying sizes, we employ different token sequence lengths while keeping a consistent patch size.
+Specifically, as depicted in Fig.
+1, for images of varying sizes, we employ different token sequence lengths while keeping a consistent patch size.
 It has been observed that the inconsistency in token length between training and prediction is a key factor in performance degradation.
@@
 Our approach demonstrates efficacy in two scenarios: zero-shot and fine-tuning.
-Extensive evaluation on datasets from five diverse tasks are conducted, including DIS5K [53], DUTS [64], ISIC [12], COD10K [16], and COCO [45].
-These datasets vary in resolution, mostly exceeding SAM’s default resolution of 1024.
+Extensive evaluations on datasets from five diverse tasks are conducted, including DIS5K [53], DUTS [64], ISIC [12], COD10K [16], and COCO [45].
+These datasets vary in resolutions, mostly exceeding SAM’s default resolution of 1024.
 In the zero-shot setting, our BA-SAM alleviates the model’s performance degradation caused by expanding the inference resolution without requiring any additional training.
@@
 Visual Foundation Models.
-Models that are trained on broad and can be adapted to numerous downstream tasks are called “Foundation Models” [5, 67, 43].
+Models that are trained on broad and can be adapted to numerous downstream tasks are called “Foundation Models” [5, 67, 43, 40, 72, 77, 84, 41].
 These models, Vision-Language Models (VLM) (CLIP [54] and DALL-E [55]) combine computer vision and natural language processing to understand and generate descriptions or analyze visual content using textual and visual information.
-Masked Image Modeling [73, 48] (MIM) refers to a technique where parts of an image are masked during training to encourage a model to learn contextual information and complete missing regions.
+Masked Image Modeling [74, 48] (MIM) masks parts of an image during the training to encourage a model to learn contextual information and complete missing regions.
 SAM [36] is a model designed for segmenting objects or areas in images, offering precise segmentation capabilities.
-In this paper, we use a variant of SAM called MobileSAM [76] as the baseline method.
-As our design optimizes the Transformer itself, it applies to all Transformer-based vision foundation models.
+We use a variant of SAM called MobileSAM [78] as the baseline method.
 Figure 2: Illustration of the proposed BA-SAM method.
 (a) In the original SAM, when input token sequences length varies during testing, the magnitude of the Softmax outputs changes drastically.
-In contrast, we propose a new scaling factor to address this issue.
-(b) We introduce a bias-mode attention mask, which produces increasing penalties on attention scores as the distance between query and key grows.
+We propose a new scaling factor to address this issue.
+(b) We introduce a bias-mode attention mask, which produces increasing penalties on attention scores as the distance between the query and key grows.
 Resolution Variation Processing.
 To enable models to be more adaptable to variations in resolutions, traditional methods to deal with VIT have relied on adjustments to positional embeddings [38] and patch sizes [4, 6, 24, 76, 44, 29, 37].
-Patch n’ Pack [13] employed sequence packing during training to handle inputs with arbitrary resolutions and aspect ratios.
+Patch n’ Pack [13] employed sequence packing during the training to handle inputs with arbitrary resolutions and aspect ratios.
 All of them necessitate training from scratch, incurring substantial computational and time costs.
-In contrast to previous approaches, we extend the concept of ’length extrapolation’ from NLP into the context of addressing scale variations in CV.
+In contrast to previous approaches, we extend the concept of length extrapolation from NLP into the context of addressing scale variations in CV.
 Length extrapolation refers to a model’s ability to generalize well to longer inputs than those it was trained on.
@@
 Parameter Efficient Tuning.
-There have been some pioneering works for the Parameter Efficient Tuning (PETL) of visual models, such as AdaptFormer [7] and visual prompt tuning (VPT) [36].
+There have been some pioneering works for the Parameter Efficient Tuning (PEFT) of visual models, such as AdaptFormer [7] and visual prompt tuning (VPT) [34].
 He et al.
@@
 They are also used for selecting paths in different branches of a network [61], or a combination of both strategies [80].
-For instance, squeeze-and-excite network [32] features an attention-like module for modeling channel-wise relationships within layer features.
-Li et al.
-[41] utilize the attention mechanism to adapt the receptive field of neurons between network branches.
 The advent of transformers has led to hybrid architectures that integrate other modules.
@@
 Segment Anything Model (SAM) [36] consists of three core modules: image encoder, prompt encoder, and mask decoder.
-The SAM model has been trained on the extensive SA-1B dataset [38], which comprises more than 1 billion automatically generated masks.
-This dataset contains 400 times more masks than any existing segmentation dataset (COCO [45], LVIS [27] and ADE20K [82]) and is accompanied by 11 million images and 1 billion masks.
+It has been trained on SA-1B dataset [36], which comprises more than 1 billion automatically generated masks.
 Consequently, SAM exhibits valuable and robust zero-shot generalization to new data without necessitating further training.
-For further details regarding the SAM methodology, readers are encouraged to refer to [38].
+further details can refer to [36].
 Our Scalable Bias-Mode Attention Mask (BA-SAM) focuses its optimization on the image encoder while keeping the structures of the mask decoder and prompt encoder unchanged.
-Scaling Factor in Attention.
+Attention in Transformer.
+In this work, we define the input sequence of image patches, $\mathbf{x}=(\mathbf{x}_{1},\dots,\mathbf{x}_{n})$ with length $N$, where $\mathbf{x}_{i}\in\mathbb{R}^{d_{x}}$.
+$q_{i}$, $k_{j}$,$v_{j}$ are calculated by $\mathbf{x}_{i}\mathbf{W}^{Q}$ , $\mathbf{x}_{j}\mathbf{W}^{K}$, $\mathbf{x}_{j}\mathbf{W}^{V}$.
+Here, the projections $\mathbf{W}^{Q},\mathbf{W}^{K},\mathbf{W}^{V}\in\mathbb{R}^{d_{x}\times d_{k}}$ are parameter matrices.
+(i) Scaling Factor.
 The two most commonly used attention functions are additive attention [1] and dot-product attention [62].
-The vanilla Transformer chooses dot-product attention because it is much faster and more space-efficient in practice.
-However, for large values of the dimension of the tokens $d_{k}$, the dot products grow large in magnitude, pushing the softmax function into regions with minimal gradients.
-They use the scaling factor $\lambda_{d}=\frac{1}{\sqrt{d_{k}}}$ to scale the dot products.
-To better analyze the role of the scaling factor, we express the output element $\mathbf{O}_{i}$ and the weight coefficient $a_{i,j}$ as follows:
-$\mathbf{O}_{i}$ $=\sum_{j=1}^{N}a_{i,j}x_{j}v_{j},\quad a_{i,j}=\frac{e^{\lambda_{d}q_{i}\cdot k_{j}}}{\sum_{j=1}^{N}e^{\lambda_{d}q_{i}\cdot k_{j}}},$ (1)
+The vanilla Transformer chooses dot-product attention for its space efficiency in practice.
+However, for larger values of $d_{k}$, the dot products grow large in magnitude, pushing the Softmax function into regions with minimal gradients.
+They use scaling factor $\lambda_{d}=\frac{1}{\sqrt{d_{k}}}$ to scale the dot products, where $d_{k}$ denotes the dimension.
+To better analyze the role of scaling factor, we express the output element $\mathbf{O}_{i}$ and the weight coefficient $a_{i,j}$ as follows:
+$\mathbf{O}_{i}$ $=\sum_{j=1}^{N}a_{i,j}v_{j},\quad a_{i,j}=\frac{e^{\lambda_{d}q_{i}\cdot k_{j}}}{\sum_{j=1}^{N}e^{\lambda_{d}q_{i}\cdot k_{j}}},$ (1)
 where $\lambda_{d}$ represents the scaling factor.
-Absolute & Relative Position Encoding in Attention.
-The original Transformer [62] incorporates absolute non-parametric positional encoding $p=(p_{1},\dots,p_{n})$ with the input token embedding $x$ as $x_{i}=x_{i}+p_{i}$, where $p_{i},x_{i}\in\mathrm{R^{d_{x}}}$.
+(ii) Absolute & Relative Position Encoding.
+The original Transformer [62] incorporates absolute non-parametric positional encoding $p=(p_{1},\dots,p_{n})$ with $x$ as $x_{i}=x_{i}+p_{i}$.
 Other works replace them with parametric encoding [23] or adopted Fourier-based kernelized versions [51].
-Absolute position encoding enforces a fixed size for all inputs.
-Recent work [56] considers the pairwise relationships between elements.
-Relative relation is important for tasks where the relative ordering or distance of the elements matters.
-This method encodes the relative position between input elements $x_{i}$ and $x_{j}$ into vectors $p_{i,j}^{v},p_{i,j}^{q},p_{i,j}^{k}\in\mathrm{R}^{d_{z}}$.
+Absolute position encoding enforces a fixed size for inputs.
+Recent work  [56] considers the pairwise relationships between elements, which encodes the relative position between input $x_{i}$ and $x_{j}$ into vectors $p_{i,j}^{v},p_{i,j}^{q},p_{i,j}^{k}\in\mathrm{R}^{d_{k}}$.
 Then, we reformulate Eq.
 (1) as follows:
-$\mathbf{O}_{i}=\sum_{j=1}^{n}a_{i,j}\left(x_{j}v_{j}+p_{i,j}^{v}\right),$ (2)
+$\mathbf{O}_{i}=\sum_{j=1}^{N}a_{i,j}\left(v_{j}+p_{i,j}^{v}\right),$ (2)
 $a_{i,j}=\frac{e^{\lambda\left(q_{i}+p_{i,j}^{q}\right)\cdot\left(k_{j}+p_{i,j}^{k}\right)}}{\sum_{j=1}^{N}e^{\lambda\left(q_{i}+p_{i,j}^{q}\right)\cdot\left(k_{j}+p_{i,j}^{k}\right)}},$ (3)
@@
 4 Methodology
-From the above section, it is evident that the original SAM’s design has a substantial limit to handling the length extrapolation problem.
-To address this, as shown in Figure 2, we present a Scalable Bias-mode Attention Mask (BA-SAM).
+Based on the preliminaries, we further analyze the characteristics of SAM: the original SAM sets the input to a fixed resolution of 1024, where it uses absolute position coding and the dot product.
+Thus, there are significant limitations in the processing of length extrapolation problems.
+To address this, as shown in Fig.
+2, we present a Scalable Bias-mode Attention Mask (BA-SAM).
 In Sec.
@@
 In Sec.
-4.2 we design a bias-mode attention mask to place more focus on neighboring tokens, ensuring the sensitivity of the model when meeting positional encoding variations.
-Finally, we explain how we will embed our proposed BA-SAM into the SAM-based structure in Sec.
+4.2 we design a bias-mode attention mask to place more focus on neighboring tokens, mitigating the impact of untrained distant information.
+Finally, we explain how we will embed our BA-SAM into the SAM-based structure in Sec.
 4.3.
 4.1 New Scaling Factor
-We discovered from the original attention module in SAM [62] that when the dot product becomes significantly large in magnitude, it can drive the Softmax layer into regions with minimal gradients.
+We observe that in the original attention module of SAM [62] when the dot product becomes significantly large in magnitude, it can drive the Softmax layer into regions with minimal gradients.
 This is because the attention distribution after Softmax becomes highly concentrated, giving rise to the issue of vanishing gradients.
 Upon closer examination of Eq.
-(1), it is obvious that the computation of the $q\cdot k$ term is intrinsically tied to both the token sequences length $N$ and the dimension $d_{k}$.
-When the token sequence length $N$ and/or the dimension $d_{k}$ significantly increases, the overall efficacy of the attention is impacted, thus leading to a noticeable performance reduction.
-To address this, we attempt to design a new scaling factor that allows the model to cope with significant variations in $N$ and $d_{k}$ values.
+(1), it is obvious that the computation of the $q\cdot k$ the term is intrinsically tied to both the token sequences length $N$ and the dimension $d_{k}$.
+When the token sequence length $N$ and the dimension $d_{k}$ significantly increase, the overall efficacy of the attention is affected, thus leading to a noticeable performance degradation.
+To address this issue, we attempt to design a new scaling factor that allows the model to cope with variations in $N$ and $d_{k}$.
 When $N$ or $d_{k}$ grows significantly, we expect to regulate the magnitude of the values within the attention layer, maintaining a similar magnitude.
-While [62] introduced a scaling factor $\lambda=\frac{1}{\sqrt{d_{k}}}$ to counteract the effect of the large growth in magnitude due to the dot products, the authors did not explain how this scaling factor was designed in the original paper.
+[62] introduced a scaling factor $\lambda=\frac{1}{\sqrt{d_{k}}}$ to counteract the effect of the large growth in magnitude due to the dot products.
 Below we will provide a theoretical derivation of this scaling factor, and then elaborate on our proposed new scaling factor.
-The dimension $d_{k}$.
-We assume that the components of $q$ and $k$ are independent Gaussian distribution variables with mean 0 and variance 1, then we want to prove that the $q\cdot k$ has mean 0 and variance $d_{k}$.
-The mean of $q\cdot k$ is:
+The dimension $d_{k}$. Following work [62], we assume that the components of $q$ and $k$ are independent random variables with mean 0 and variance 1. The mean of $q\cdot k$ is:
 $\mathrm{E}[q\cdot k]$ $=\mathrm{E}\left[\sum_{i=1}^{d_{k}}q_{i}k_{i}\right]=\sum_{i=1}^{d_{k}}\mathrm{E}\left[q_{i}\right]\mathrm{E}\left[k_{i}\right]=0$ (4)
@@
 Given this, we can approximately consider the $q\cdot k$ values to be within the range of $-3\sqrt{d_{k}}$ to $3\sqrt{d_{k}}$, according to properties of Gaussian distribution.
-For larger models, $d_{k}$ is generally a larger positive value, resulting in a significant increase in the magnitude of numerical values of $q\cdot k$, compared to the additive attention option which has the range of $\left[-3,3\right]$.
+For larger models, $d_{k}$ is generally a larger positive value, resulting in a significant increase in the magnitude of numerical values of $q\cdot k$, compared to the additive attention option, which has the range of $\left[-3,3\right]$.
 Consequently, the attention distribution after Softmax becomes highly concentrated.
-This leads to severe gradient vanishing, which in turn hampers the effectiveness of training and can lead to less desired performance.
+This leads to severe gradient vanishing, which hampers the effectiveness of the training and can lead to less desired performance.
 As the $q\cdot k$ values lie in the range of $[-3\sqrt{d_{k}},3\sqrt{d_{k}}]$, the scaling factor can be simply defined as $\lambda_{d}=\frac{1}{\sqrt{d_{k}}}$, in order to maintain a similar magnitude.
-Our new scaling factor.
-We provided the interpretation on how the original scaling factor was designed, which has been overlooked by the original authors.
-Now we explain the design of our new scaling factor.
-Our goal is to regulate the magnitude of the Softmax value $a_{ij}$ (Eq.
-(1)) as the token sequences length $N$ and the dimension $d_{k}$ change.
-We further denote the token sequences length during the training as $N_{train}$ and the token sequence length during the testing as $N_{test}$, where $N_{test}>>N_{train}$.
-We now consider how to maintain a similar magnitude of $a_{ij}$ during training and testing times.
-We introduce the scaling factor $\lambda_{N}$, which is related to $N$ and controls the output scale of the Softmax value $a_{ij}$ , namely, $S(a_{i,j}^{N_{test}})=S(a_{i,j}^{N_{train}})$.
-Here, $S()$ represents the scale estimation function.
-We next substitute Eq.
-(1) into $S(a_{i,j}^{N_{test}})=S(a_{i,j}^{N_{train}})$, and simplify $q_{i}\cdot k_{j}$ into $x_{i,j}$ provided the magnitude of $q\cdot k$ has been already maintained.
-Now we have:
-$S(\frac{e^{\lambda_{n}x_{i,j}}}{\sum_{j=1}^{n}e^{\lambda_{n}x_{i,j}}})=S(\frac{e^{x_{i,j}}}{\sum_{j=1}^{m}e^{x_{i,j}}})$ (6)
-$\Rightarrow S({e^{\lambda_{n}}})=S(\frac{N_{test}}{N_{train}})\Rightarrow\lambda_{n}=log_{N_{train}}N_{test}$
+Our new scaling factor. We have provided the interpretation of how the original scaling factor was designed. Now, we explain the design of our new scaling factor.
+According to Eq.
+4 and Eq.
+5, the scale of $q\cdot k$ has been consistent with the additive attention by $\lambda_{d}$, which can be seen as $a_{i,j}$ is independent of $d_{k}$.
+We simplify $\lambda_{d}q_{i}\cdot k_{j}$ into $x_{i,j}$ and further discuss the effect of length $N$ on $a_{i,j}$.
+In Eq.1 , $a_{i,j}$ can be seen as the conditional distribution with $i$ as the condition and $j$ as the random variable.
+Inspired by [60], we introduce information entropy to constrain $a_{i,j}$.
+Specifically, entropy is a measure of uncertainty, and we expect the uncertainty of $a_{i,j}$ to be insensitive to the length $N$ (i.e., the value of each $a_{i,j}$ will change when the token increases, but the entropy value of the overall $a_{i,j}$ can remain relatively stable).
+The entropy of $a_{i,j}$ is :$\mathcal{H}_{i}=-\sum_{j=1}^{N}a_{i,j}\log a_{i,j}$ and we substitute Eq.
+(1):
+$\mathcal{H}_{i}=\log\sum_{j=1}^{N}e^{\lambda_{n}x_{i,j}}-\frac{\sum_{j=1}^{N}e^{\lambda_{n}x_{i,j}}\left(\lambda_{n}x_{i,j}\right)}{\sum_{j=1}^{N}e^{\lambda_{n}x_{i,j}}}$ (6)
+Then, we substitute the approximate estimates into Eq. (6):
+$\sum_{j=1}^{N}e^{x_{i,j}}=N\times\frac{1}{N}\sum_{j=1}^{N}e^{x_{i,j}}\approx N\mathbb{E}_{j}\left[e^{x_{i,j}}\right]$ (7)
+$\mathbb{E}_{j}\left[e^{x_{i,j}}\left(x_{i,j}\right)\right]\approx 0,\mathbb{E}_{j}\left[e^{x_{i,j}}\right]=O(1)$
+We wish to use $\lambda_{n}$ to offset the effect of $N$ on $\mathcal{H}_{i}$. Then, we have the following result:
+$\mathcal{H}_{i}\approx\log N-k\lambda_{n}=0\Rightarrow\lambda_{n}=\frac{\log N}{k}$ (8)
+where $k$ is a parameter value.
+We denote the token sequence length during the training as $N_{train}$ and the token sequence length during the testing as $N_{test}$, where $N_{test}>>N_{train}$.
+When $N=N_{train}$, $\lambda_{n}=1$ (consistent with the training length).
+As such, $k=\log N_{train}$ and finally we have $\lambda_{n}=log_{N_{train}}N_{test}$.
 Considering both $\lambda_{d}$ and $\lambda_{n}$, we can ultimately derive our new scaling factor as:
 $\lambda=\lambda_{d}\lambda_{n}=\frac{log_{N_{train}}N_{test}}{\sqrt{d_{k}}}$ (9)
-Our new scaling factor in Eq.
-(7) ensures the magnitude of values during attention computation remains consistent, regardless of variations in $d_{k}$ and $N$.
-It will enhance the extrapolative capacity of the model.
+Our new scaling factor in Eq. (9) ensures attention computation remains consistent, regardless of variations in $d_{k}$ and $N$. It will enhance the extrapolative capacity of the model.
 4.2 Bias-Mode Attention Mask
 Another challenge is that changes in token sequence length will lead to variations in positional encoding.
-It is important to ensure the insensitivity of the model when such positional encoding variations occur during testing.
+It is important to ensure the insensitivity of the model when such positional encoding variations occur during the testing.
 One possible way is absolute encoding without trainable parameters, such as Sinusoidal [62].
 It requires the position encoding to have strong local-to-global inference capabilities.
-As mentioned by Su [60], this resembles Taylor series approximations: by knowing the values of several derivatives at a particular point, effective estimates can be made within a neighborhood.
 Nevertheless, this assumes that the given function has high-order smoothness (higher-order derivatives exist and are bounded).
-Commonly used positional encodings are often combined with trigonometric functions, essentially high-frequency oscillatory functions with periodicity.
-However, these methods fail to satisfy the requirement of bounded high-order derivatives, making it less accurate to estimate the extrapolated results.
-Another potential approach is using local attention, which constrains the model’s field of view and remains insensitive to variations in token sequence length.
+Commonly-used positional encodings are often combined with trigonometric functions These methods fail to satisfy the requirement of bounded high-order derivatives, making it less accurate to estimate the extrapolated results.
+Another potential approach is using local attention [47], which constrains the model’s field of view and remains insensitive to variations in token sequence length.
 However, local attention is typically implemented using a local window, necessitating modifications to the SAM structure, which requires re-training from scratch and is unacceptable for deployment in the downstream tasks.
-To this end, we propose to enable the attention layer to put more focus on neighboring tokens of the current token.
+To this end, we propose enabling the attention layer to focus more on the current token’s neighboring tokens.
 As such, even with an increase in the length of a token sequence, each token is scarcely affected by the untrained tokens from distant positions.
@@
 We introduce a bias after the query-key dot product.
-As shown in Figure 3, this mask exhibits a bias specified on the distance between the query-key pairs (i.e., $q\cdot k$).
+As shown in Fig.
+3, this mask exhibits a bias specified on the distance between the query-key pairs (i.e., $q\cdot k$).
 We expect that this proposed mask imposes penalties on attention scores between distant query-key pairs, and the penalty increases as the distance between a key $q$ and a query $k$ grows.
@@
 We further discuss the setting of $\beta$ based on different cases.
-When conducting zero-shot transfer learning without fine-tuning, we set $\beta$ to a static, non-learned fixed value.
-The specific value setting will be discussed in the experimental section (Sec.
+When conducting zero-shot generalization without fine-tuning, we set $\beta$ to a static, non-learned fixed value.
+The experimental section will discuss the specific value setting (Sec.
 5).
 When fine-tuning is required, we make $\beta$ trainable.
-Since our Bias-Mode Attention Mask is lightweight relative to the model structure, it incurs negligible training time overhead.
+Since our Bias-Mode Attention Mask is lightweight relative to the model structure, it incurs negligible training cost overhead.
 Figure 3: Embedding of our BA-SAM into a SAM backbone. NSF indicates our new scaling factor, and BM-AM denotes our designed bias-mode attention mask.
@@
 4.3 BA-SAM Model
-As shown in Figure 3, our BA-SAM is simple to implement, and can be seamlessly integrated into SAM [38] and its variants.
+As shown in Fig.
+3, our BA-SAM is simple to implement, and can be seamlessly integrated into SAM [36] and its variants.
 Specifically, our design involves a new scaling factor (NSF) for the attention layer and a bias-mode attention mask (BM-AM).
@@
 With fine-tuning
-MobileSAM [76] 1024 1024 0.098 - 0.034 - 0.057 - 0.030 -
-2048 0.520 42.2% 0.082 4.8% 0.065 0.8% 0.063 3.3%
-4096 0.542 44.4% 0.235 20.1% 0.101 4.4% 0.138 10.8%
+MobileSAM [78] 1024 1024 0.059 - 0.034 - 0.057 - 0.030 -
+2048 0.144 8.5 % 0.082 4.8% 0.065 0.8% 0.063 3.3%
+4096 0.205 14.6 % 0.235 20.1% 0.101 4.4% 0.138 10.8%
 2048 2048 0.083 - 0.045 - 0.056 - 0.036 -
 4096 0.227 14.4% 0.091 4.6% 0.066 1.0% 0.059 2.3%
-Ours (w  [76]) 1024 1024 0.037 - 0.031 - 0.055 - 0.029 -
-2048 0.499 40.4% 0.078 4.4% 0.061 0.6% 0.057 2.5%
-4096 0.531 43.6% 0.226 19.2% 0.098 4.3% 0.118 8.6%
+Ours (w  [78]) 1024 1024 0.053 - 0.031 - 0.055 - 0.029 -
+2048 0.118 6.5 % 0.078 4.4% 0.061 0.6% 0.057 2.5%
+4096 0.190 13.7% 0.226 19.2% 0.098 4.3% 0.118 8.6%
 2048 2048 0.080 - 0.043 - 0.053 - 0.033 -
@@
 We employed the widely-used MAE (Mean Absolute Error) score.
-Lower MAE scores indicate better model accuracy.
+Lower MAE scores indicate better model performance.
 $\Delta$diff denotes the performance degradation due to resolution changes.
-Compared to the SAM [38] and MobileSAM [76] baselines, our BA-SAM achieves smaller degradation when encountering token sequences length changes.
+Compared to the SAM [36] and MobileSAM [78] baselines, our proposed BA-SAM achieves smaller degradation when encountering token sequence length changes.
 The best MAE performance is highlighted in bold, and the smallest performance degradation is underlined.
-We first describe the datasets and implementation details in Sec.
-5.1.
-Then, we compare BA-SAM’s performance with state-of-the-art methods across multiple tasks in Sec.
-5.2.
-Finally, extensive ablation studies and analysis are conducted in Sec.
-5.3.
-to demonstrate the effectiveness.
 5.1 Datasets and Implementations
@@
 For a comprehensive evaluation of BA-SAM, we conduct extensive experiments on a wide range of segmentation tasks, i.e., salient object segmentation [18, 15, 20], complex object segmentation [19], skin lesion segmentation [96], camouflaged object detection [17], which correspond to four datasets: DUTS [64], DIS-TE4 [53], ISIC [12] and COD10K [16].
-Besides, we also verify it on the challenging COCO [45] instance segmentation benchmark.
-More details about datasets can be referred to in the supplementary material.
+Besides, we verify it on the challenging COCO [45] instance segmentation benchmark.
+More details are referred to in the supplementary material.
 Implementation Details.
@@
 More details are provided in the supplementary material.
-The code will be released at this link.
 Evaluation metrics.
@@
 5.2 Results
-Results of Various Object Segmentation Tasks: Table 1 demonstrates the effectiveness of our approach across four diverse segmentation datasets.
+Results of Various Object Segmentation Tasks: Tab.
+1 demonstrates the effectiveness of our approach across four diverse segmentation datasets.
 $\Delta$diff denotes the value of the performance degradation due to resolution changes during the inference.
-The upper and lower parts of the table indicate the results without and with fine-tuning, respectively due to input resolution variations.
+The upper and lower parts of the table indicate the results without and with fine-tuning.
 The best MAE performance is highlighted in bold, and the smallest degradation is underlined.
@@
 5.3 for details.
-Results of Instance Segmentation: In Table 3, we evaluate the performance of our method on the COCO [45] instance segmentation benchmark.
+Results of Instance Segmentation: In Tab.
+3, we evaluate the performance of our method on the COCO [45] instance segmentation benchmark.
 For a fair comparison, all experiments are conducted in a zero-shot manner, with the same initialization parameters for the comparative methods and without the use of any additional training data.
 Our BA-SAM consistently outperforms SAM [36] and MobileSAM [78] baselines, demonstrating better zero-shot generalization capability on instance segmentation.
-Comparisons with State-of-the-Art Methods: To further demonstrate the superiority of the effectiveness and generalizability of our method, we compare the state-of-the-art approaches in Table 2.
+Comparisons with State-of-the-Art Methods: To further demonstrate the superiority of the effectiveness and generalizability of our method, we compare the state-of-the-art approaches in Tab.
+2.
 From the table, we have two following observations: Firstly, all the state-of-the-art approaches [57, 79, 46, 95, 83, 53, 35, 16] show less-desirable performances in each dataset.
@@
 Importantly, unlike [46, 57, 79, 12], we eliminate the need for employing additional techniques to further enhance the performance.
-As shown in Table 2, our generalized model also consistently promotes the performance of SAM on all datasets, demonstrating its remarkable generalizability.
+As shown in Tab.
+2, our generalized model also consistently promotes the performance of SAM on all datasets, demonstrating its remarkable generalizability.
 5.3 Ablation Study and Analysis
@@
 Ablations Studies of Each Component.
-Table 4 summarizes the effect of each designed component on the settings with and without fine-tuning, respectively.
+Tab.
+4 summarizes the effect of each designed component on the settings with and without fine-tuning, respectively.
 The baseline means using the MobileSAM [78] as the base network that uses the vanilla scaling factor (VSF) in the attention layer [62].
@@
 From the table, we observe that NSF could achieve superior performances than the baseline with VSF.
-This is because the vanilla attention in SAM [38] and MobileSAM [76] does not consider maintaining the magnitude consistency of the Softmax outputs changes drastically when input resolutions vary during testing.
+This is because the vanilla attention in SAM [36] and MobileSAM [78] does not consider maintaining the magnitude consistency when Softmax outputs change drastically due to input resolutions vary during the testing.
 In contrast, our NSF explicitly maintains the magnitude consistency and alleviates the performance degradation.
@@
 The same framework, 4 Specialized models
-Ours (BA-SAM) 0.037 0.031 0.055 0.029
+Ours (BA-SAM) 0.053 0.031 0.055 0.029
 Generalized model
@@
 SAM [36] 42.5 69.6 44.7 29.7 47.0 56.7
-Ours (w [38]) 43.0 70.0 45.4 30.0 47.6 57.4
+Ours (w [36]) 43.0 70.0 45.4 30.0 47.4 57.1
 MobileSAM [78] 40.8 68.4 41.6 26.0 44.4 57.6
-Ours (w [76]) 41.6 69.0 42.3 26.2 44.8 58.4
+Ours (w [78]) 41.2 69.0 42.1 26.2 44.8 58.2
 Table 3: Results (AP) on COCO [45] instance segmentation.
@@
 In the Bias-Mode Attention Mask, the magnitude of the slope $\beta$ determines penalty rates in different heads.
-In the zero-shot case, we use a fixed slope.
-Table 5 reveals the impact of different slopes of BM-AM on the ISIC [12] dataset.
 We found that the best performance is achieved when $\beta=0.1$.
 Besides, our method is robust to different slope choices.
-In addition to the zero-shot setting, we also validate the BM-AM in fine-tuning scenarios on four datasets, shown in Table 1.
-The learnable parameters brought more performance growth compared with the fixed slope, while also bringing a slight increase in computation.
+In the zero-shot case, we use a fixed slope $\beta=1$ by default.
+(More details could be referred to in the supplementary material)
 Computational Efficiency.
-In Table 6, we analyze the computational efficiency between the baselines and our BA-SAM.
+In Tab.
+5, we analyze the computational efficiency between the baselines and our BA-SAM.
 All the experiments are conducted on the same NVIDIA RTX 4090GPU to ensure fair comparisons.
@@
 Table 4: Ablation study of each component on the settings with and without fine-tuning. Numbers indicate the performance degradation, $\Delta$diff. A lower $\Delta$diff means a better performance.
-Size 0.1 0.5 1 1.5 2 2.5
-MAE 0.0375 0.0383 0.0381 0.0393 0.0388 0.0391
-$\Delta$diff 40.4 40.8 40.4 40.9 41.2 40.7
-Table 5: Impact of the slope $\beta$ in BM-AM (zero-shot setting).
 Model Params (M) Speed (ms) Train Hours (h)
@@
 In this paper, we address the important problem of varying image resolutions in SAM models by reformulating it as a problem of length extrapolation.
-To enhance the length extrapolation capability of SAM, we propose Scalable Bias-mode Attention Mask (BA-SAM).
+To enhance the length extrapolation capability of SAM, we propose the Scalable Bias-mode Attention Mask (BA-SAM) SAM.
 A new scaling factor (NSF) is introduced to maintain the consistent magnitude of attention.
 In addition, a bias-mode attention mask (BM-AM) is designed to prioritize neighboring information, mitigating the impact of untrained distant information.
-Extensive evaluation on diverse datasets reveals its ability to significantly alleviate performance degradation in the zero-shot setting and achieve state-of-the-art performance with minimal fine-tuning.
-Furthermore, we proposed a generalized model and benchmark, showcasing BA-SAM’s generalizability across all four datasets simultaneously.
+Extensive evaluation of diverse datasets reveals its ability to significantly alleviate performance degradation in the zero-shot setting and achieve state-of-the-art performance with minimal fine-tuning.
+Furthermore, we proposed a generalized model and benchmark, showcasing BA-SAM’s generalizability across all four datasets.
+7 Acknowledgement
+Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102), Shanghai Science and Technology Commission (21511101200), National Natural Science Foundation of China (No.
+72192821), YuCaiKe [2023] Project Number: 14105167-2023